### v1

# April 3, 2025

Abstract.
In this paper, we establish the existence of solutions for time-dependent linearly viscoelastic bodies confined to a specified half-space. The confinement condition under consideration is of Signorini type, and is given over the boundary of the linearly viscoelastic body under consideration. In order to derive a suitable notion of solution that possesses a sufficient degree of regularity, we introduce the concept of “admissible” applied body force. In agreement with expectations, the variational problem that solutions must satisfy is a system of hyperbolic variational inequalities, and the displacement acceleration is a vector-valued measure in the sense of Dinculeanu.
Keywords. Variational Inequalities $\cdot$ Penalty method $\cdot$ Asymptotic Analysis $\cdot$ Obstacle Problems
MSC 2020. 35L85, 47H07, 74B10.
1. Introduction
The contact of elastic and viscoelastic materials with rigid obstacles arises in many applicative fields, including biomedical engineering, aerospace design, and materials science. For instance, the motion of aortic heart valves - modelled as linearly elastic shells constrained within a spatial domain to avoid collisions - has inspired mathematical frameworks for obstacle problems in elasticity [22, 26, 39]. Similarly, the confinement of deformable structures like polymer films or viscoelastic membranes in prescribed half-spaces is critical for applications ranging from soft robotics to biomechanical implants [38].
The mechanical deformation of linearly viscoelastic bodies is classically modelled using energy functionals from linearized elasticity theory [7, 22]. Recent advances extend these principles to dynamic and time-dependent settings. For example, Bock and Jarušek analysed unilateral dynamic contact problems for viscoelastic Reissner-Mindlin plates [2], while [3] established existence results for thermoelastic von Kármán plates vibrating against rigid obstacles. A key challenge in such problems lies in reconciling hyperbolic variational inequalities for displacements with memory-dependent constitutive laws, as seen in quasi-static viscoelasticity with self-contact [28, 42]. The latter introduces novel solution concepts distinct from classical frameworks [4, 30], emphasising admissible forces that preclude terminal-phase collisions, a paradigm central to our work as well.
Recent studies highlight unresolved questions in viscoelastic obstacle problems, such as the interplay between memory kernel decay rates and solution regularity [12]. These align with broader inquiries into non-local integro-differential equations governing viscoelastic dynamics, where semigroup methods and spectral analysis remain pivotal [12]. Meanwhile, confinement conditions in linearized elasticity, particularly for three-dimensional bodies in half-spaces, demand sophisticated Sobolev-space analyses to ensure that suitable energy estimates hold [15].
In this paper, we formulate a three-dimensional Signorini problem for a time-dependent linearly viscoelastic body constrained to a half-space. To the best of our knowledge, the current literature does not address the study of time-dependent obstacle problems for three-dimensional viscoelastic bodies subjected to a confinement condition Signorini-type. Our approach synthesizes ideas from [42] - where forces appear to avoid terminal-phase contact - and the geometric confinement frameworks of [11]. The main result of this paper, namely Theorem 4.4, establishes existence of solutions under “admissible” applied body forces. The concept of solution for the governing model is recovered by two layer of approximations: in the first layer, we relax the constraint by penalising the classical energy in linearised viscoelasticity, while in the second layer we discretise the relaxed problem by Galerkin method. A noticeable feature that we establish is that the displacement acceleration enjoys the regularity of a vector-valued measure in the sense of Dinculeanu [13] or, equivalently, the displacement velocity is a function of bounded variation [29].
The paper is divided into four sections, including this one. In section 2 we present the main notation from elasticity and differential geometry. In section 3 we formulate the Signorini problem for a three-dimensional viscoelastic body. Finally, in section 4, we recover the existence of solutions for the model under consideration by means of compactness methods.
2. Geometrical preliminaries
For details about the classical notions of differential geometry recalled in this section, see, e.g. [8] or [9].
Latin indices, except when they are used for indexing sequences, take their values in the set $\{1,2,3\}$, and the summation convention with respect to repeated indices is systematically used in conjunction with this rule.
For comprehensive information on the classical concepts of differential geometry referenced in this section, see, e.g., [8] or [9].
Latin indices, except when utilized for indexing sequences, take values from the set ${1,2,3}$, and the summation convention for repeated indices is consistently applied alongside this rule.
For an open subset $\Omega\subset\mathbb{R}^{3}$, the notations $L^{2}(\Omega)$ and $H^{1}(\Omega)$ refer to the standard Lebesgue and Sobolev spaces, respectively. The notation $\mathcal{D}(\Omega)$ indicates the space of functions that are infinitely differentiable on $\Omega$ and have compact support within $\Omega$. The notation $|\cdot|_{X}$ denotes the norm in a vector space $X$. Spaces of vector-valued functions are represented by boldface letters. Lebesgue-Bochner spaces (see, e.g., [29]) are represented by the notation $L^{p}(0,T;X)$, where $1\leq p\leq\infty$, $T>0$ and $X$ is a Banach space satisfying the Radon-Nikodym property. The notation $X^{\ast}$ refers to the dual space of a vector space $X$, and the notation $\langle\cdot,\cdot\rangle_{X^{\ast},X}$ denotes the duality pairing between $X^{\ast}$ and $X$. The notation $\mathcal{C}^{0}([0,T];X)$ designates the space of continuous $X$-valued functions defined over the compact interval $[0,T]$, while the special notation $\langle\langle\cdot,\cdot\rangle\rangle_{X^{\ast},X}$ denotes the duality pair between $(\mathcal{C}^{0}([0,T];X))^{\ast}$ and $\mathcal{C}^{0}([0,T];X)$. The notation $\mathcal{M}([0,T];X^{\ast})$ denotes the space of $X^{\ast}$-valued measures defined over the compact interval $[0,T]$.
[…]
A Lipschitz domain $\Omega\subset\mathbb{R}^{3}$ is a non-empty, open, bounded, and connected subset with a Lipschitz continuous boundary $\Gamma$, where the set $\Omega$ is locally on the same side of $\Gamma$. The notation $\,\mathrm{d}x$ indicates the volume element in $\Omega$, and the symbol $\,\mathrm{d}\Gamma$ indicates the area element along $\Gamma$. Finally, let $\Gamma=\Gamma_{0}\cup\Gamma_{1}$ be a $\,\mathrm{d}\Gamma$-measurable portion of the boundary such that $\Gamma_{0}\cap\Gamma_{1}=\emptyset$ and area $\textup{area }\Gamma_{0}>0$. We denote by $\nu$ the outer-unit normal vector field along the boundary $\Gamma$ of the set $\Omega$.
As a model of the three-dimensional “physical” space $\mathbb{R}^{3}$, we consider a real three-dimensional affine Euclidean space, which is defined by selecting a point $O$ as the origin and associating it with a real three-dimensional Euclidean space, denoted $\mathbb{E}^{3}$. We equip $\mathbb{E}^{3}$ with an orthonormal basis consisting of three vectors $\bm{e}^{i}$. The Euclidean inner product of two elements $\bm{a}$ and $\bm{b}$ in $\mathbb{E}^{3}$ is denoted by $\bm{a}\cdot\bm{b}$; the Euclidean norm of any $\bm{a}\in\mathbb{E}^{3}$ is denoted by $|\bm{a}|$; the Kronecker symbol is denoted by $\delta^{ij}$.
The characterization of $\mathbb{R}^{3}$ as an affine Euclidean space implies that with each point $x\in\mathbb{R}^{3}$ is associated a uniquely defined vector $\bm{Ox}\in\mathbb{E}^{3}$. The origin $O\in\mathbb{R}^{3}$ and the orthonormal vectors $\bm{e}^{i}\in\mathbb{E}^{3}$ together form a Cartesian frame in $\mathbb{R}^{3}$, and the three components $x_{i}$ of the vector $\bm{Ox}$ relative to the basis $\bm{e}^{i}$ are referred to as the Cartesian coordinates of $x\in\mathbb{R}^{3}$, or the Cartesian components of $\bm{Ox}\in\mathbb{E}^{3}$. Once a Cartesian frame has been established, any point $x\in\mathbb{R}^{3}$ can be identified with the vector $\bm{Ox}=x_{i}\bm{e}^{i}\in\mathbb{E}^{3}$. We then denote $\partial_{i}=\partial/\partial x_{i}$.
The set $\overline{\Omega}$ is the reference configuration occupied by a linearly elastic body in the absence of applied body forces. We assume that $\overline{\Omega}$ is in a natural state, meaning the body is stress-free in this configuration. Following [7], we also assume that the material is isotropic, homogeneous, and linearly elastic. Under these conditions, the behaviour of the linearly elastic material is governed by its two Lamé constants $\lambda\geq 0$ and $\mu>0$. We denote the viscosity constants by $\theta\geq 0$ and $\xi>0$. The positive constant $\rho$ designates the mass density of the linearly viscoelastic body per unit volume.
In what follows, “a.e.” stands for “almost everywhere” and “a.a.” stands for “almost all”. Define the space
$\bm{V}(\Omega):=\{\bm{v}=(v_{i})\in\bm{H}^{1}(\Omega);\bm{v}={\bf{0}}\textup{ on }\Gamma_{0}\},$
and equip it with the norm
$\|\bm{v}\|_{\bm{V}(\Omega)}:=\left(\sum_{i}\|v_{i}\|_{H^{1}(\Omega)}^{2}\right)^{1/2}.$
We define the three-dimensional elasticity tensor in Cartesian coordinates and we denote its components by $A^{ijk\ell}$. We recall that the contravariant components of this tensor are defined by (see, e.g., [7])
$A^{ijk\ell}:=\lambda\delta^{ij}\delta^{k\ell}+\mu(\delta^{ik}\delta^{j\ell}+\delta^{i\ell}\delta^{jk}),$
and that $A^{ijk\ell}=A^{jik\ell}=A^{k\ell ij}\in\mathcal{C}^{1}(\overline{\Omega})$.
We then define the three-dimensional viscoelasticity tensor in Cartesian coordinates and we denote its components by $B^{ijk\ell}$. We recall that the contravariant components of this tensor are defined by (see, e.g., [22])
$B^{ijk\ell}:=\theta\delta^{ij}\delta^{k\ell}+\xi(\delta^{ik}\delta^{j\ell}+\delta^{i\ell}\delta^{jk}),$
and that $B^{ijk\ell}=B^{jik\ell}=B^{k\ell ij}\in\mathcal{C}^{1}(\overline{\Omega})$.
For each $\bm{v}\in\bm{H}^{1}(\Omega)$ we consider the linearised change of metric tensor $\bm{e}(\bm{v})$, whose components $e_{i\|j}(\bm{v})$ are defined by:
$e_{i\|j}(\bm{v}):=\dfrac{1}{2}(\partial_{j}v_{i}+\partial_{i}v_{j})\in L^{2}(\Omega).$
This tensor is symmetric, i.e., $e_{i\|j}(\bm{v})=e_{j\|i}(\bm{v})$, for all $\bm{v}\in\bm{H}^{1}(\Omega)$.
We now state Korn’s inequality in Cartesian coordinates (see, e.g., Theorem 6.3-6 of [7]).
Theorem 2.1.
Let $\Omega$ be a Lipschitz domain in $\mathbb{R}^{3}$ and let $\Gamma_{0}$ be a non-zero area subset of the whole boundary $\Gamma$. Then, there exists a constant $c_{0}>0$ such that
$c_{0}^{-1}\|\bm{v}\|_{\bm{H}^{1}(\Omega)}\leq\|\bm{e}(\bm{v})\|_{\bm{L}^{2}(\Omega)}\leq c_{0}\|\bm{v}\|_{\bm{H}^{1}(\Omega)},$
for all $\bm{v}\in\bm{V}(\Omega)$. ∎
Several proofs have been provided for this intricate inequality. Notably, see [17], [20], [24], [25], page 110 of [14], and Section 6.3 of [32]. In [41], Korn’s inequality is demonstrated in the space $W^{1,p}(\Omega)$ for $1<p\leq\infty$. An elementary proof can be found in [33], along with additional details in Appendix (A) of [31].
We denote by $\mathcal{A}:\bm{V}(\Omega)\to\bm{V}^{\ast}(\Omega)$ the linear and continuous operator defined in a way that:
$\langle\mathcal{A}\bm{w},\bm{v}\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}:=\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{w})e_{i\|j}(\bm{v})\,\mathrm{d}x,\quad\textup{ for all }\bm{w},\bm{v}\in\bm{V}(\Omega).$
Thanks to the properties of the fourth order three-dimensional elasticity tensor $\{A^{ijk\ell}\}$ and Korn’s inequality in Cartesian coordinates (Theorem 2.1), we have that the operator $\mathcal{A}$ is symmetric and uniformly positive-definite, in the sense that there exists a constant $C_{e}^{(1)}>0$ such that:
$\sum_{i,j}|t_{ij}|^{2}\leq C_{e}^{(1)}A^{ijk\ell}(x)t_{k\ell}t_{ij},$
for all $x\in\overline{\Omega}$ and all symmetric matrices $(t_{ij})$.
In a completely similar fashion, we denote by $\mathcal{B}:\bm{V}(\Omega)\to\bm{V}^{\ast}(\Omega)$ the linear and continuous operator defined in a way that:
$\langle\mathcal{B}\bm{w},\bm{v}\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}:=\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\bm{w})e_{i\|j}(\bm{v})\,\mathrm{d}x,\quad\textup{ for all }\bm{w},\bm{v}\in\bm{V}(\Omega).$
Thanks to the properties of the fourth order three-dimensional elasticity tensor $\{B^{ijk\ell}\}$ and Korn’s inequality in Cartesian coordinates (Theorem 2.1), we have that the operator $\mathcal{B}$ is symmetric and uniformly positive-definite (cf., e.g., [22]), in the sense that there exists a constant $C_{e}^{(2)}>0$ such that:
$\sum_{i,j}|t_{ij}|^{2}\leq C_{e}^{(2)}B^{ijk\ell}(x)t_{k\ell}t_{ij},$
for all $x\in\overline{\Omega}$ and all symmetric matrices $(t_{ij})$.
3. Variational formulation of the time-dependent obstacle problem for a linearly viscoelastic body
In this paper, we examine a specific obstacle problem for a linearly elastic body subjected to a confinement condition. This condition requires that any admissible displacement vector field $v_{i}\bm{e}^{i}$ must ensure that all points of the corresponding deformed configuration remain within a half-space of the form
:=$\mathbb{H}:=\{x\in\mathbb{R}^{3};\bm{Ox}\cdot\bm{q}\geq 0\},$
where $\bm{q}$ is a unit-norm vector given once and for all. Let us denote by $\bm{I}$ the identity mapping $\bm{I}:\overline{\Omega}\to\mathbb{E}^{3}$ and let us assume that the undeformed reference configuration satisfies
$\bm{I}(x)\cdot\bm{q}\geq 0,\quad\textup{ for all }x\in\overline{\Omega},$
or, in other words, there may be contact between the obstacle and the reference configuration when no applied forces are acting on the reference configuration.
The general confinement condition can be thus formulated by requiring that any admissible displacement vector field must satisfy
$(\bm{I}(x)+v_{i}(x)\bm{e}^{i})\cdot\bm{q}\geq 0,$
for all $x\in\Gamma$ or, possibly, only for a.a. $x\in\Gamma$ when the covariant components $v_{i}$ are required to belong to the space $H^{1}(\Omega)$. The latter is a confinement condition of Signorini type (cf., e.g., [27]).
Let us introduce the subset $\bm{U}(\Omega)$ of admissible displacements takes the following form:
$\bm{U}(\Omega):=\{\bm{v}=(v^{i})\in\bm{V}(\Omega);(\bm{I}+v_{i}\bm{e}^{i})\cdot\bm{q}\geq 0\,\textup{a.e. on }\Gamma\}.$
The linearly viscoelastic body under consideration is subjected to both applied body forces, that we denote by $\bm{f}=(f_{i})\in L^{2}(0,T;\bm{L}^{2}(\Omega))$. The applied surface forces for the problem under consideration are expressed in terms of the vectors $\bm{e}^{i}$ of the Cartesian framework. For the sake of simplicity, we do not consider the action of surface forces, even though it is licit to consider the action of traction forces on the portion of the boundary that does not engage contact with the obstacle.
Moreover, we restrict ourselves to considering applied body forces in the sense of the definition stated below.
Definition 3.1.
We say that an applied body force $\bm{f}=(f_{i})\in L^{2}(0,T;\bm{L}^{2}(\Omega))$ is admissible if there exists $T_{0}=T_{0}(\bm{f})\in(0,T)$ such that the constraint becomes inactive for a.a. $T_{0}<t<T$.
By restricting ourselves to considering applied body forces that are admissible in the sense of Definition 3.1, we are allowed to incorporate this information in the governing model for the problem under consideration, and we thus solely need to focus on making sure that the constraint is implemented for a.a. $0<t<T_{0}$. This specialisation of the applied body forces will be critical for establishing the sought existence result. In connection with this assumption, we mention the recent paper [42] where the authors considered a variational formulation of a collision problem non-linearly viscoelastic solids. The concept of solution proposed in [42] suggests that the applied force shuts down and that the terminal point $T$ is so large so as to have zero velocity in terminal part of the motion. The case considered in [42] is a particular case of “admissible” applied body force in the sense of Definition 3.1.
In the problem here discussed, we propose a different concept of solution that is more similar to the classical concept of solutions for variational inequalities [6, 19, 30]. In our formulation, the velocity of the linearly viscoelastic body needs not vanishing in the long run; however, we need to incorporate in the governing model the requirement that the constraint is inactive in the terminal part of the motion.
The variational problem $\mathcal{P}(\Omega)$ indicated below constitutes the point of arrival of our analysis. The aim of this paper is to show that one such problem admits a solution enjoying the regularity announced below.
In all what follows, we restrict ourselves to considering applied body forces that are “admissible” in the sense of Definition 3.1.
Problem $\mathcal{P}(\Omega)$.
Find $\bm{u}=(u_{i}):(0,T)\to\bm{V}(\Omega)$ such that
$\bm{u}\in L^{\infty}(0,T;\bm{U}(\Omega)),$
$\dot{\bm{u}}\in L^{\infty}(0,T;\bm{L}^{2}(\Omega))\cap L^{2}(0,T;\bm{V}(\Omega)),$
$\ddot{\bm{u}}\in\mathcal{M}([0,T];\bm{V}^{\ast}(\Omega)),$
that satisfies the following variational inequalities
$2\rho\langle\langle\ddot{\bm{u}},\bm{v}\rangle\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}-2\rho\langle\dot{\bm{u}}(T),\bm{u}(T)\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}+2\rho\int_{\Omega}\bm{u}_{1}\cdot\bm{u}_{0}\,\mathrm{d}x+2\rho\|\dot{\bm{u}}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{2}$
$\quad+\int_{0}^{T}\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}(t))e_{i\|j}(\bm{v}(t)-\bm{u}(t))\,\mathrm{d}x\,\mathrm{d}t+\int_{0}^{T}\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\dot{\bm{u}}(t))e_{i\|j}(\bm{v}(t))\,\mathrm{d}x\,\mathrm{d}t$
$\quad-\dfrac{1}{2}\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\bm{u}(T))e_{i\|j}(\bm{u}(T))\,\mathrm{d}x\,\mathrm{d}t+\dfrac{1}{2}\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\bm{u}_{0})e_{i\|j}(\bm{u}_{0})\,\mathrm{d}x$
$\geq\int_{0}^{T}\int_{\Omega}f^{i}(t)(v_{i}(t)-u_{i}(t))\,\mathrm{d}x\,\mathrm{d}t,$
for all $\bm{v}=(v_{i})\in\mathcal{C}^{0}([0,T];\bm{U}(\Omega))$, and that satisfies the following initial conditions
$\begin{cases}\bm{u}(0)=\bm{u}_{0},\\ \dot{\bm{u}}(0)=\bm{u}_{1},\end{cases}$
where $\bm{u}_{0}=(u_{i,0})\in\bm{U}(\Omega)$, and $\bm{u}_{1}=(u_{i,1})\in\bm{V}(\Omega)$ are prescribed. $\blacksquare$
4. Existence of solutions for Problem $\mathcal{P}(\Omega)$
The formulation of this initial boundary value problem is inspired by the classical Signorini formulation (cf., e.g., [27]). We construct solutions of Problem $\mathcal{P}(\Omega)$ by studying its penalised version first. In what follows, we denoted by $\kappa>0$ a penalty parameter that is meant to approach zero. In what follows, we denote by $\{f\}^{-}$ the negative part of a function $f$, which is defined as $\{f\}^{-}:=-\min\{f,0\}$.
Problem $\mathcal{P}_{\kappa}(\Omega)$.
Find $\bm{u}_{\kappa}=(u_{i,\kappa}):(0,T)\to\bm{V}(\Omega)$ such that
$\bm{u}_{\kappa}\in L^{\infty}(0,T;\bm{V}(\Omega)),$
$\dot{\bm{u}}_{\kappa}\in L^{\infty}(0,T;\bm{L}^{2}(\Omega))\cap L^{2}(0,T;\bm{V}(\Omega)),$
$\ddot{\bm{u}}_{\kappa}\in L^{\infty}(0,T;\bm{V}^{\ast}(\Omega)),$
that satisfies the following variational equations
$2\rho\,\int_{\Omega}\ddot{\bm{u}}_{\kappa}(t)\cdot\bm{v}\,\mathrm{d}x+\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}_{\kappa}(t))e_{i\|j}(\bm{v})\,\mathrm{d}x\,\mathrm{d}t$
$\quad+\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\dot{\bm{u}}_{\kappa}(t))e_{i\|j}(\bm{v})\,\mathrm{d}x\,\mathrm{d}t$
$\quad-\dfrac{\chi_{(0,T_{0})}(t)}{\kappa}\int_{\Gamma}\{[\bm{I}+u_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\}^{-}v_{i}\bm{e}^{i}\cdot\bm{q}\,\mathrm{d}\Gamma$
$=\int_{\Omega}f^{i}(t)v_{i}\,\mathrm{d}x\,\mathrm{d}t,$
for all $\bm{v}=(v_{i})\in\bm{V}(\Omega)$ in the sense of distributions in $(0,T)$, and that satisfies the following initial conditions
$\begin{cases}\bm{u}_{\kappa}(0)=\bm{u}_{0},\\ \dot{\bm{u}}_{\kappa}(0)=\bm{u}_{1},\end{cases}$
where $\bm{u}_{0}=(u_{i,0})\in\bm{U}(\Omega)$, and $\bm{u}_{1}=(u_{i,1})\in\bm{V}(\Omega)$ are the same as in Problem $\mathcal{P}(\Omega)$. $\blacksquare$
Penalised problems akin to Problem $\mathcal{P}_{\kappa}(\Omega)$ have been addressed in the literature by several authors; see for instance [1, 2, 3]. In [3], the solution strategy exploits the underlying regularity of the model as well as the fact that the model is posed over a two-dimensional Lipschitz domain. In [1, 2], the argument hinges on the fact that the competitors to the role of solutions are chosen among those whose weak derivative in time exists and is of class $L^{2}$. The competitors to the role of solutions for Problem $\mathcal{P}(\Omega)$ are instead only of class $\mathcal{C}^{0}$ in time.
In the case treated in this paper, for a.a. $t\in(0,T)$, the displacement solving Problem $\mathcal{P}(\Omega)$ is, in general, of class $\bm{H}^{1}(\Omega)$, where $\Omega$ is a Lipschitz domain in $\mathbb{R}^{3}$.
Let us recall the following general result which can be found in, e.g., Lemma 3.1 in [36] or [16].
Lemma 4.1.
Let $\mathscr{O}\subset\mathbb{R}^{m}$, with $m\geq 1$ an integer, be an open set. The operator {}$-\{\cdot\}^{-}:L^{2}(\mathscr{O})\to L^{2}(\mathscr{O})$ defined by
$f\in L^{2}(\mathscr{O})\mapsto-\{f\}^{-}:=\min\{f,0\}\in L^{2}(\mathscr{O}),$
is monotone, bounded and Lipschitz continuous with Lipschitz constant equal to $1$. ∎
We re-cast the penalty term, which describes the energy that has to be paid by the system to violate the constraint, in terms of the non-linear operator $\mathcal{N}:\bm{V}(\Omega)\to\bm{V}^{\ast}(\Omega)$, that we define by:
$\langle\mathcal{N}\bm{w},\bm{v}\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}:=-\int_{\Gamma}\{[\bm{I}+w_{i}\bm{e}^{i}]\cdot\bm{q}\}^{-}v_{i}\bm{e}^{i}\cdot\bm{q}\,\mathrm{d}\Gamma,$
for all $\bm{w}=(w_{i})\in\bm{V}(\Omega)$ and all $\bm{v}=(v_{i})\in\bm{V}(\Omega)$, and we observe that the latter definition makes sense being the trace of an element in $\bm{V}(\Omega)$ of class $\bm{L}^{2}(\Gamma)$ when $\Omega$ is a Lipschitz domain as in our case. As a result, it is straightforward to observe that also the operator $\mathcal{N}$ is monotone, bounded and Lipschitz continuous.
We then define the non-linear operator $\tilde{\mathcal{N}}:L^{2}(0,T;\bm{V}(\Omega))\to L^{2}(0,T;\bm{V}^{\ast}(\Omega))$ pointwise by
(1) $(\tilde{\mathcal{N}}\bm{w})(t):={\mathcal{N}}(\bm{w}(t)),\quad\textup{ for a.a. }t\in(0,T),$
and we observe that, in light of the properties of the non-linear operator $\mathcal{N}$ introduced beforehand, the operator $\tilde{\mathcal{N}}$ is monotone, bounded and Lipschitz continuous.
We now recall the very important Gronwall inequality often used to study evolutionary problems: (see, e.g. the original paper [21] or, for instance, Theorem 1.1 in Chapter III of [23]).
Theorem 4.1.
Let $T>0$ and suppose that the function $y:[0,T]\to\mathbb{R}$ is absolutely continuous and such that
$\dfrac{\,\mathrm{d}y}{\,\mathrm{d}t}(t)\leq a(t)y(t)+b(t),\quad\textup{ for a.a. }t\in(0,T),$
where $a,b\in L^{1}(0,T)$ and $a(t),b(t)\geq 0$ for a.a. $t\in(0,T)$. Then, it results
$y(t)\leq\left[y(0)+\int_{0}^{t}b(s)\,\mathrm{d}s\right]\exp\left(\int_{0}^{t}a(s)\,\mathrm{d}s\right),\quad\textup{ for all }t\in[0,T].$
Theorem 4.2.
Assume that the applied body force $\bm{f}=(f_{i})\in L^{2}(0,T;\bm{L}^{2}(\Omega))$ is admissible in the sense of Definition 3.1. Then, Problem $\mathcal{P}_{\kappa}(\Omega)$ admits at least one solution.
Proof.
For the sake of clarity, we break the proof into three parts, numbered (i)–(iii).
(i) Construction of a Galerkin approximation. Since $\bm{V}(\Omega)$ is an infinite dimensional separable Hilbert space which is also dense in $\bm{L}^{2}(\Omega)$, we infer that there exists an orthogonal basis $\{\bm{w}^{p}\}_{p=1}^{\infty}$ of the space $\bm{V}(\Omega)$, whose elements also constitute a Hilbert basis of the space $\bm{L}^{2}(\Omega)$.
The existence of such a basis is assured by the spectral theorem (Theorem 6.2-1 of [37]). For each positive integer $m\geq 1$, we denote by $\bm{E}^{m}$ the following $m$-dimensional linear hull:
$\bm{E}^{m}:=\textup{Span }\{\bm{w}^{p}\}_{p=1}^{m}\subset\bm{V}(\Omega)\subset\bm{L}^{2}(\Omega).$
Since each element of this Hilbert basis is independent of the variable $t$, we have that $\bm{w}^{p}\in L^{\infty}(0,T;\bm{V}(\Omega))$ for each integer $1\leq p\leq m$. We now discretise Problem $\mathcal{P}_{\kappa}(\Omega)$.
Problem $\mathcal{P}_{\kappa}^{m}(\Omega)$.
Find functions $c_{p,\kappa}:[0,T]\to\mathbb{R}$, $1\leq p\leq m$, such that
$\bm{u}^{m}_{\kappa}(t):=\sum_{p=1}^{m}c_{p,\kappa}(t)\bm{w}^{p},\quad\textup{ for a.a. }t\in(0,T),$
and satisfying the following variational equations
$2\rho\int_{\Omega}\ddot{\bm{u}}_{\kappa}^{m}(t)\cdot\bm{w}^{p}\,\mathrm{d}x+\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}_{\kappa}^{m}(t))e_{i\|j}(\bm{w}^{p})\,\mathrm{d}x\,\mathrm{d}t$
$\quad+\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\dot{\bm{u}}_{\kappa}^{m}(t))e_{i\|j}(\bm{w}^{p})\,\mathrm{d}x\,\mathrm{d}t$
$\quad-\dfrac{\chi_{(0,T_{0})}(t)}{\kappa}\int_{\Gamma}\{[\bm{I}+u_{i,\kappa}^{m}(t)\bm{e}^{i}]\cdot\bm{q}\}^{-}w_{i}^{p}\bm{e}^{i}\cdot\bm{q}\,\mathrm{d}\Gamma$
$=\int_{\Omega}f^{i}(t)w_{i}^{p}\,\mathrm{d}x\,\mathrm{d}t,$
for all $1\leq p\leq m$ in the sense of distributions in $(0,T)$, and that satisfies the following initial conditions
$\begin{cases}\bm{u}_{\kappa}^{(m)}(0)=\bm{u}_{0}^{m},\\ \dot{\bm{u}}_{\kappa}^{(m)}(0)=\bm{u}_{1}^{m},\end{cases}$
where the initial conditions $\bm{u}_{0}^{m}$ and $\bm{u}_{1}^{m}$ are defined by:
$\bm{u}_{0}^{m}:$ $=\sum_{p=1}^{m}\int_{\Omega}\left\{(\bm{u}_{0}\cdot\bm{w}^{p})\bm{w}^{p}+(\partial_{i}\bm{u}_{0}\cdot\partial_{i}\bm{w}^{p})\bm{w}^{p}\right\}\,\mathrm{d}x,$
$\bm{u}_{1}^{m}:$ $=\sum_{p=1}^{m}\int_{\Omega}(\bm{u}_{1}\cdot\bm{w}^{p})\bm{w}^{p}\,\mathrm{d}x.$
$\blacksquare$
We observe that $\bm{u}_{0}^{m}\to\bm{u}_{0}$ in $\bm{V}(\Omega)$ as $m\to\infty$, and that $\bm{u}_{1}^{m}\to\bm{u}_{1}$ in $\bm{L}^{2}(\Omega)$ as $m\to\infty$. Since the coefficients $c_{p,\kappa}$ and their derivatives only depend on the time variable, we can take them outside of the integral sign, getting a $m\times m$ non-linear system of second order ordinary differential equations with respect to the variable $t$. Such a system can be rewritten in the form
(2) $2\rho\ddot{\bm{C}}_{\kappa}(t)$ $=\left(-\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{w}^{r})e_{i\|j}(\bm{w}^{p})\,\mathrm{d}x\right)_{p,r=1}^{m}\bm{C}_{\kappa}(t)$
$\quad+\left(-\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\bm{w}^{r})e_{i\|j}(\bm{w}^{p})\,\mathrm{d}x\right)_{p,r=1}^{m}\dot{\bm{C}}_{\kappa}(t)$
$\quad+\dfrac{\chi_{(0,T_{0})}(t)}{\kappa}\left(\int_{\Gamma}\left(\{[\bm{I}+(\bm{C}_{\kappa}(t)\cdot(w^{1}_{i}\dots w^{m}_{i}))\bm{e}^{i}]\cdot\bm{q}\}^{-}\right)(w^{p}_{i}\bm{e}^{i}\cdot\bm{q})\,\mathrm{d}x\right)_{p=1}^{m}$
$\quad+\left(\int_{\Omega}f^{i}(t)w_{i}^{p}\,\mathrm{d}x\right)_{p=1}^{m}$
where $\bm{C}_{\kappa}(t):=(c_{1,\kappa}(t)\dots c_{m,\kappa}(t))$. Thanks to Lemma 4.1, the right hand side of (2) is Lipschitz continuous in $\mathbb{R}^{m}$ uniformly with respect to $t$, since it does not explicitly depend on $t$. An application of Theorem 1.45 of [40] gives that for each integer $m\geq 1$ there exists a unique global solution $\bm{u}_{\kappa}^{m}$ to Problem $\mathcal{P}_{\kappa}^{m}(\Omega)$, defined a.e. over the interval $(0,T)$, such that:
$\bm{u}^{m}_{\kappa}\in L^{\infty}(0,T;\bm{E}^{m}),$
$\dot{\bm{u}}^{m}_{\kappa}\in L^{\infty}(0,T;\bm{E}^{m}),$
$\ddot{\bm{u}}^{m}_{\kappa}\in L^{\infty}(0,T;\bm{E}^{m}).$
(ii) Energy estimates for the approximate solutions. Let us multiply the variational equations in Problem $\mathcal{P}_{\kappa}^{m}(\Omega)$ by $\dot{c}_{p,\kappa}(t)$, with $0<t<T$, and sum with respect to $p\in\{1,\dots,m\}$. The penalised variational equations in Problem $\mathcal{P}_{\kappa}^{m}(\Omega)$ take the form
(3) $\rho\dfrac{\,\mathrm{d}}{\,\mathrm{d}t}\int_{\Omega}\dot{u}_{i,\kappa}^{m}(t)\dot{u}_{i}^{m}(t)\,\mathrm{d}x+\dfrac{1}{2}\dfrac{\,\mathrm{d}}{\,\mathrm{d}t}\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}^{m}_{\kappa}(t))e_{i\|j}(\bm{u}^{m}_{\kappa}(t))\,\mathrm{d}x$
$\quad+\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\dot{\bm{u}}^{m}_{\kappa}(t))e_{i\|j}(\dot{\bm{u}}^{m}_{\kappa}(t))\,\mathrm{d}x$
$\quad+\dfrac{\chi_{(0,T_{0})}(t)}{2\kappa}\dfrac{\,\mathrm{d}}{\,\mathrm{d}t}\left(\int_{\Gamma}\left(\left\{[\bm{I}+u^{m}_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\right)^{2}\,\mathrm{d}x\right)$
$=\int_{\Omega}f^{i}(t)\dot{u}^{m}_{i,\kappa}(t)\,\mathrm{d}x,$
and are valid in the sense of distributions in $(0,T)$. Observe that the differentiation of the negative part is obtained as a result of the same computational steps as in Stampacchia’s theorem (cf., e.g., Theorem 4.4 on page 153 of [16]), together with an application of Theorem 8.28 of [29]. An integration over the interval $(0,t)$, where $0<t\leq T$, changes (3) into:
(4) $\rho\int_{\Omega}\dot{u}_{i,\kappa}^{m}(t)\dot{u}_{i,\kappa}^{m}(t)\,\mathrm{d}x+\dfrac{1}{2}\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}^{m}_{\kappa}(t))e_{i\|j}(\bm{u}^{m}_{\kappa}(t))\,\mathrm{d}x$
$\quad+\int_{0}^{t}\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\dot{\bm{u}}^{m}_{\kappa}(t))e_{i\|j}(\dot{\bm{u}}^{m}_{\kappa}(t))\,\mathrm{d}x\,\mathrm{d}t$
$\quad+\dfrac{\chi_{(0,T_{0})}(t)}{2\kappa}\int_{\Gamma}\left(\left\{[\bm{I}+u^{m}_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\right)^{2}\,\mathrm{d}x$
$=\rho\int_{\Omega}u_{i,1}^{m}u_{i,1}^{m}\,\mathrm{d}x+\dfrac{1}{2}\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}_{0}^{m})e_{i\|j}(\bm{u}_{0}^{m})\,\mathrm{d}x$
$\quad+\dfrac{\chi_{(0,T_{0})}(t)}{2\kappa}\int_{\Gamma}\left(\left\{[\bm{I}+u^{m}_{i,0}\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\right)^{2}\,\mathrm{d}x$
$\quad+\int_{0}^{t}\int_{\Omega}f^{i}(\tau)\dot{u}^{m}_{i,\kappa}(\tau)\,\mathrm{d}x\,\mathrm{d}\tau.$
Let us observe that the third integral in the left hand side of (4) tends to zero, since $\bm{u}_{0}^{m}\to\bm{u}_{0}$ in $\bm{V}(\Omega)$ as $m\to\infty$, and $\bm{u}_{0}\in\bm{U}(\Omega)$. An application of the Cauchy-Schwarz inequality gives that:
(5) $\int_{0}^{t}\int_{\Omega}f^{i}(\tau)\dot{u}^{m}_{i,\kappa}(\tau)\,\mathrm{d}x\,\mathrm{d}\tau$
$\leq\left(\int_{0}^{T}\|\bm{f}(t)\|_{\bm{L}^{2}(\Omega)}^{2}\,\mathrm{d}t\right)^{1/2}\left(\int_{0}^{t}\|\dot{\bm{u}}^{m}_{\kappa}(\tau)\|_{\bm{L}^{2}(\Omega)}^{2}\,\mathrm{d}\tau\right)^{1/2}$
$\leq\dfrac{c_{0}C_{e}^{(2)}}{2}\int_{0}^{T}\|\bm{f}(t)\|_{\bm{L}^{2}(\Omega)}^{2}\,\mathrm{d}t+\dfrac{1}{2c_{0}C_{e}^{(2)}}\int_{0}^{t}\|\dot{\bm{u}}^{m}_{\kappa}(\tau)\|_{\bm{L}^{2}(\Omega)}^{2}\,\mathrm{d}\tau.$
By the uniform positive-definiteness of the elasticity tensors $\{A^{ijk\ell}\}$ and $\{B^{ijk\ell}\}$, Korn’s inequality (Theorem 2.1), (4), and (5), the following estimate holds for $m$ sufficiently large:
(6) $\rho\|\dot{\bm{u}}^{m}_{\kappa}(t)\|_{\bm{L}^{2}(\Omega)}^{2}+\dfrac{\|\bm{u}^{m}_{\kappa}(t)\|_{\bm{V}(\Omega)}^{2}}{2c_{0}C_{e}^{(1)}}+\dfrac{1}{2c_{0}C_{e}^{(2)}}\int_{0}^{t}\|\dot{\bm{u}}^{m}_{\kappa}(\tau)\|_{\bm{V}(\Omega)}^{2}\,\mathrm{d}\tau+\dfrac{\chi_{(0,T_{0})}(t)}{2\kappa}\left\|\left\{[\bm{I}+u^{m}_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\right\|_{L^{2}(\Gamma)}^{2}$
$\leq\rho\|\bm{u}_{1}\|_{\bm{L}^{2}(\Omega)}^{2}+\dfrac{c_{0}C_{e}^{(2)}}{2}\|\bm{f}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{2}+\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}_{0}^{m})e_{i\|j}(\bm{u}_{0}^{m})\,\mathrm{d}x$
$\quad+\rho\int_{0}^{t}\|\dot{\bm{u}}^{m}_{\kappa}(\tau)\|_{\bm{L}^{2}(\Omega)}^{2}\,\mathrm{d}\tau+\dfrac{1}{2c_{0}C_{e}^{(1)}}\int_{0}^{t}\|\bm{u}^{m}_{\kappa}(\tau)\|_{\bm{V}(\Omega)}^{2}\,\mathrm{d}\tau$
$\quad+\dfrac{1}{2\kappa}\int_{0}^{t}\chi_{(0,T_{0})}(\tau)\left\|\left\{[\bm{I}+u^{m}_{i,\kappa}(\tau)\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\right\|_{L^{2}(\Gamma)}^{2}\,\mathrm{d}\tau.$
An application of the Gronwall’s inequality (Theorem 4.1) with
$y(t):=\rho\int_{0}^{t}\|\dot{\bm{u}}^{m}_{\kappa}(\tau)\|_{\bm{L}^{2}(\Omega)}^{2}\,\mathrm{d}\tau+\dfrac{1}{2c_{0}C_{e}^{(1)}}\int_{0}^{t}\|\bm{u}^{m}_{\kappa}(\tau)\|_{\bm{V}(\Omega)}^{2}\,\mathrm{d}\tau+\dfrac{1}{2\kappa}\int_{0}^{t}\chi_{(0,T_{0})}(\tau)\left\|\left\{[\bm{I}+u^{m}_{i,\kappa}(\tau)\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\right\|_{L^{2}(\Gamma)}^{2}\,\mathrm{d}\tau,$
$a\equiv 1>0$ and
$b\equiv\max\left\{1,\rho,\dfrac{c_{0}C_{e}^{(2)}}{2}\right\}\left(\|\bm{u}_{1}\|_{\bm{L}^{2}(\Omega)}^{2}+\|\bm{f}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{2}+\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}_{0})e_{i\|j}(\bm{u}_{0})\,\mathrm{d}x\right)\geq 0,$
gives the following uniform upper bound
(7) $\rho\int_{0}^{t}\|\dot{\bm{u}}^{m}_{\kappa}(\tau)\|_{\bm{L}^{2}(\Omega)}^{2}\,\mathrm{d}\tau+\dfrac{1}{2c_{0}C_{e}^{(1)}}\int_{0}^{t}\|\bm{u}^{m}_{\kappa}(\tau)\|_{\bm{V}(\Omega)}^{2}\,\mathrm{d}\tau$
$\quad+\dfrac{1}{2\kappa}\int_{0}^{t}\chi_{(0,T_{0})}(\tau)\left\|\left\{[\bm{I}+u^{m}_{i,\kappa}(\tau)\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\right\|_{L^{2}(\Gamma)}^{2}\,\mathrm{d}\tau$
$\leq T\max\left\{1,\rho,\dfrac{c_{0}C_{e}^{(2)}}{2}\right\}\left(\|\bm{u}_{1}\|_{\bm{L}^{2}(\Omega)}^{2}+\|\bm{f}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{2}+\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}_{0})e_{i\|j}(\bm{u}_{0})\,\mathrm{d}x\right)e^{T},$
for all $t\in[0,T]$. Therefore, we obtain that
[…]
and, moreover, by (6), there exists a positive uniform constant $L$ such that
(9) $0\leq\left\|\chi_{(0,T_{0})}\left\{[\bm{I}+u^{m}_{i,\kappa}\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\right\|_{L^{2}(0,T;L^{2}(\Gamma))}\leq\sqrt{\kappa Tbe^{T}}.$
Since the following direct sum decomposition holds
$\bm{V}(\Omega)=\bm{E}^{m}\oplus(\bm{E}^{m})^{\perp},$
we get that for any $\bm{v}\in\bm{V}(\Omega)$ with $\|\bm{v}\|_{\bm{V}(\Omega)}=1$, and a.a. $t\in(0,T)$, the variational equations in Problem $\mathcal{P}_{\kappa}^{m}(\Omega)$ give
$|\langle\ddot{\bm{u}}^{m}_{\kappa}(t),\bm{v}\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}|\leq\|\bm{f}(t)\|_{\bm{L}^{2}(\Omega)}+\left(\max_{\overline{\Omega}}|A^{ijk\ell}|\right)\|\bm{u}^{m}_{\kappa}\|_{L^{\infty}(0,T;\bm{V}(\Omega))}$
$\quad+\left(\max_{\overline{\Omega}}|B^{ijk\ell}|\right)\|\dot{\bm{u}}^{m}_{\kappa}(t)\|_{\bm{V}(\Omega)}+\dfrac{1}{\kappa}\left\|\left\{[\bm{I}+u^{m}_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\bm{q}\right\|_{\bm{L}^{2}(\Gamma)},$
and, by (8) and (9), we thus infer that there exists a constant $C>0$, independent of $m$, $t$ and $\kappa$, such that:
(10) $\|\ddot{\bm{u}}^{m}_{\kappa}\|_{L^{2}(0,T;\bm{V}^{\ast}(\Omega))}\leq C\left(1+\dfrac{1}{\kappa}\right).$
In light of the assumed admissibility of the applied body force $\bm{f}$ (Definition 3.1), we have that for a.a. $T_{0}<t<T$, there exists a constant $C=C(T_{0})>0$, independent of $m$, $t$ and $\kappa$, such that:
(11) $\|\ddot{\bm{u}}^{m}_{\kappa}\|_{L^{2}(T_{0},T;\bm{V}^{\ast}(\Omega))}\leq C.$
(iii) Passage to the limit as $m\to\infty$ and completion of the proof. By (8), (9), and (10) we can infer that there exist subsequences, still denoted $\{\bm{u}^{m}\}_{m=1}^{\infty}$, $\{\dot{\bm{u}}^{m}_{\kappa}\}_{m=1}^{\infty}$ and $\{\ddot{\bm{u}}^{m}_{\kappa}\}_{m=1}^{\infty}$ such that the following convergences hold:
[…]
By the Sobolev embedding theorem (cf., e.g., Theorem 10.1.20 of [34]), we obtain
(13) $\bm{u}^{m}_{\kappa}\rightharpoonup\bm{u}_{\kappa},\quad\textup{ in }\mathcal{C}^{0}([0,T];\bm{V}(\Omega))\textup{ as }m\to\infty,$
$\dot{\bm{u}}^{m}_{\kappa}\rightharpoonup\dot{\bm{u}}_{\kappa},\quad\textup{ in }\mathcal{C}^{0}([0,T];\bm{V}^{\ast}(\Omega))\textup{ as }m\to\infty,$
An application of Theorem 8.28 of [29] to the fourth convergence of the process (12) gives:
(14) $\kappa^{-1}\chi_{(0,T_{0})}\left\{[\bm{I}+u^{m}_{i,\kappa}\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\rightharpoonup\chi_{\kappa},\quad\textup{ in }L^{2}((0,T)\times\Omega)\textup{ as }m\to\infty.$
Thanks to Lemma 4.1, the first convergence of (13) and the weak convergence (14), Theorem 8.28 of [29], Theorem 8.62 of [29] and the monotonicity of the operator $\tilde{\mathcal{N}}$ defined in (1), we are in a position to apply Theorem 9.13-2 of [10] and, so, to obtain:
(15) $\chi_{\kappa}=\kappa^{-1}\chi_{(0,T_{0})}\left\{[\bm{I}+u_{i,\kappa}\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\in L^{2}((0,T)\times\Omega).$
We now verify that $\bm{u}_{\kappa}$ is a solution of the variational equations in Problem $\mathcal{P}_{\kappa}^{m}(\Omega)$. Let $\psi\in\mathcal{D}(0,T)$ and let $\tilde{m}\geq 1$ be any integer. For each $m\geq\tilde{m}$, we have
(16) $2\rho\int_{0}^{T}\int_{\Omega}\ddot{u}_{i,\kappa}^{m}(t)v_{i}\,\mathrm{d}x\psi(t)\,\mathrm{d}t+\int_{0}^{T}\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}^{m}_{\kappa}(t))e_{i\|j}(\bm{v})\,\mathrm{d}x\psi(t)\,\mathrm{d}t$
$\quad+\int_{0}^{T}\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\dot{\bm{u}}^{m}_{\kappa}(t))e_{i\|j}(\bm{v})\,\mathrm{d}x\psi(t)\,\mathrm{d}t$
$\quad-\dfrac{1}{\kappa}\int_{0}^{T}\chi_{(0,T_{0})}(t)\int_{\Omega}\left(\left\{[\bm{I}+u^{m}_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\right\}^{-}\right)(v_{i}\bm{e}^{i}\cdot\bm{q})\,\mathrm{d}x\psi(t)\,\mathrm{d}t$
$=\int_{0}^{T}\int_{\Omega}f^{i}(t)v_{i}\,\mathrm{d}x\psi(t)\,\mathrm{d}t,$
for all $\bm{v}\in\bm{E}^{\tilde{m}}$. Thanks to Lemma 4.1, the convergence process (12), (15), the arbitrariness of $\psi\in\mathcal{D}(0,T)$, as well as the fact that
$\overline{\bigcup_{\tilde{m}\geq 1}\bm{E}^{\tilde{m}}}^{\|\cdot\|_{\bm{V}(\Omega)}}=\bm{V}(\Omega),$
a passage to the limit as $m\to\infty$ in (16) shows that $\bm{u}_{\kappa}$ is a solution to the penalised variational equations in Problem $\mathcal{P}_{\kappa}(\Omega)$.
The last property that we have to check is the validity of the initial conditions for $\bm{u}_{\kappa}$. Let us introduce the operator $\bm{L}_{0}:\mathcal{C}^{0}([0,T];\bm{V}(\Omega))\to\bm{V}(\Omega)$ defined in a way such that $\bm{L}_{0}(\bm{v}):=\bm{v}(0)$. Such an operator $\bm{L}_{0}$ turns out to be linear and continuous and, therefore, by the first convergence of (13), we immediately obtain:
$\bm{u}_{\kappa}(0)=\bm{u}_{0},\quad\textup{ in }\bm{L}^{2}(\Omega).$
Similarly, let us introduce the operator $\bm{L}_{1}:\mathcal{C}^{0}([0,T];\bm{V}^{\ast}(\Omega))\to\bm{V}^{\ast}(\Omega)$ defined in a way such that $\bm{L}_{1}(\bm{v}):=\bm{v}(0)$. Such an operator $\bm{L}_{1}$ turns out to be linear and continuous and since $\bm{u}^{m}_{1}$ is the projection of $\bm{u}_{1}$ onto $\bm{E}^{m}$, the second convergence of (13) gives:
$\bm{u}^{m}_{1}\rightharpoonup\dot{\bm{u}}_{\kappa}(0)=\bm{u}_{1},\quad\textup{in }\bm{V}^{\ast}(\Omega)\textup{ as }m\to\infty.$
We thus obtain that $\bm{u}_{\kappa}$ is a solution of Problem $\mathcal{P}_{\kappa}^{m}(\Omega)$, and the proof is complete. ∎
The next step consists in showing that the sequence of solutions $\{\bm{u}_{\kappa}\}_{\kappa>0}$ of Problem $\mathcal{P}_{\kappa}(\Omega)$ converges in some suitable sense to a solution of Problem $\mathcal{P}(\Omega)$. Before establishing this result, that constitutes the main contribution of this paper, we need to establish some preparatory lemmas.
Lemma 4.2.
The operator $\mathcal{A}:\bm{V}(\Omega)\to\bm{V}^{\ast}(\Omega)$ defined by
$\langle\mathcal{A}\bm{w},\bm{v}\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}:=\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{w})e_{i\|j}(\bm{v})\,\mathrm{d}x,$
is linear and continuous. Moreover, also the corresponding time-dependent version $\tilde{\mathcal{A}}:L^{\infty}(0,T;\bm{V}(\Omega))\to L^{\infty}(0,T;\bm{V}^{\ast}(\Omega))$ defined pointwise by
$(\tilde{\mathcal{A}}\bm{w})(t):=\mathcal{A}(\bm{w}(t)),\quad\textup{ for all }\bm{w}\in L^{\infty}(0,T;\bm{V}(\Omega)),$
for a.a. $t\in(0,T)$ is linear and continuous.
Proof.
The operator $\mathcal{A}$ is the classical operator in linearised elasticity; its linearity and boundedness are straightforward to establish (cf., e.g., [7]). For what concerns the time-dependent version of this operator, the linearity and boundedness follow, in the same spirit as in [35], from the uniform boundedness of the components $A^{ijk\ell}$ of the fourth order three-dimensional elasticity tensor. ∎
In a similar fashion, the following preparatory result can be established.
Lemma 4.3.
The operator $\mathcal{B}:\bm{V}(\Omega)\to\bm{V}^{\ast}(\Omega)$ defined by
$\langle\mathcal{B}\bm{w},\bm{v}\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}:=\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\bm{w})e_{i\|j}(\bm{v})\,\mathrm{d}x,$
is linear and continuous. Moreover, also the corresponding time-dependent version ~$\tilde{\mathcal{B}}:L^{2}(0,T;\bm{V}(\Omega))\to L^{2}(0,T;\bm{V}^{\ast}(\Omega))$ defined pointwise by
(~$(\tilde{\mathcal{B}}\bm{w})(t):=\mathcal{B}(\bm{w}(t)),\quad\textup{ for all }\bm{w}\in L^{2}(0,T;\bm{V}(\Omega)),$
for a.a. $t\in(0,T)$ is linear and continuous. ∎
Lemma 4.4.
Consider the sequence $\{\bm{u}_{\kappa}\}_{\kappa>0}$, where each $\bm{u}_{\kappa}$ is a solution of Problem $\mathcal{P}_{\kappa}(\Omega)$. Up to passing to a subsequence, the following convergences hold:
[…]
$\dot{\bm{u}}_{\kappa}$ $\rightharpoonup\dot{\bm{u}},\quad\textup{ in }L^{2}(0,T;\bm{V}(\Omega))\textup{ as }\kappa\to 0,$
$\ddot{\bm{u}}_{\kappa}$ $\rightharpoonup\ddot{\bm{u}},\quad\textup{ in }L^{2}(T_{0},T;\bm{V}^{\ast}(\Omega))\textup{ as }\kappa\to 0,$
$\chi_{(0,T_{0})}\{[\bm{I}+u_{i,\kappa}\bm{e}^{i}]\cdot\bm{q}\}^{-}$ $\to\chi_{(0,T_{0})}\{[\bm{I}+u_{i}\bm{e}^{i}]\cdot\bm{q}\}=0,\quad\textup{ in }L^{2}(0,T;\bm{L}^{2}(\Gamma))\textup{ as }\kappa\to 0,$
$\bm{u}_{\kappa}$ $\rightharpoonup\bm{u},\quad\textup{ in }\mathcal{C}^{0}([0,T];\bm{V}(\Omega))\textup{ as }\kappa\to 0,$
$\dot{\bm{u}}_{\kappa}(T)$ $\rightharpoonup\dot{\bm{u}}(T),\quad\textup{ in }\bm{V}^{\ast}(\Omega)\textup{ as }\kappa\to 0.$
In particular, we obtain that $\bm{u}\in\mathcal{C}^{0}([0,T];\bm{U}(\Omega))$ and $\bm{u}(0)=\bm{u}_{0}\in\bm{U}(\Omega)$.
Proof.
Since the right-hand sides of the estimates in (7), (9) and (11) are independent of $m$ and $\kappa$, we obtain that the following convergences hold up to passing to a subsequence:
[…]
$\dot{\bm{u}}_{\kappa}$ $\rightharpoonup\dot{\bm{u}},\quad\textup{ in }L^{2}(0,T;\bm{V}(\Omega))\textup{ as }\kappa\to 0,$
[…]
The convergence $\bm{u}_{\kappa}\rightharpoonup\bm{u}$ in $\mathcal{C}^{0}([0,T];\bm{L}^{2}(\Omega))$ as $\kappa\to 0$ holds thanks to Theorem 10.1.20 in [34] (see also [18]). The weak convergences of the sequences $\{\dot{\bm{u}}_{\kappa}\}_{\kappa>0}$ and $\{\ddot{\bm{u}}_{\kappa}\}_{\kappa>0}$ and the assumed admissibility of the applied body force $\bm{f}$ in turn imply that:
$\dot{\bm{u}}_{\kappa}(T)\rightharpoonup\dot{\bm{u}}(T),\quad\textup{ in }\bm{V}^{\ast}(\Omega)\textup{ as }\kappa\to 0.$
Let us now observe that an application of (9) gives
$\int_{0}^{T}\chi_{(0,T_{0})}\langle\mathcal{N}\bm{u}_{\kappa}(t),\bm{u}_{\kappa}(t)\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\,\mathrm{d}t$
$=\int_{0}^{T}\int_{\Gamma}\chi_{(0,T_{0})}\left|\{[\bm{I}+u_{i,\kappa}\bm{e}^{i}]\cdot\bm{q}\}^{-}\right|^{2}\,\mathrm{d}\Gamma\,\mathrm{d}t+\int_{0}^{T}\int_{\Gamma}\chi_{(0,T_{0})}\{[\bm{I}+u_{i,\kappa}\bm{e}^{i}]\cdot\bm{q}\}^{-}(\bm{I}\cdot\bm{q})\,\mathrm{d}\Gamma\,\mathrm{d}t\to 0,$
as $\kappa\to 0$. The Lipschitz continuity and monotonicity of the operator $\tilde{\mathcal{N}}$ introduced in (1) put us in position to apply Theorem 9.13-2 in [10] so as to infer that
$\{[\bm{I}+u_{i}\bm{e}^{i}]\cdot\bm{q}\}^{-}=0,\quad\textup{ in }L^{2}(T_{0},T;\bm{V}^{\ast}(\Omega)),$
as well as the validity of the fifth convergence in (17). In particular, the latter implies that $\bm{u}\in\mathcal{C}^{0}([0,T];\bm{U}(\Omega))$. Clearly, the assumed admissibility of the applied body forces (Definition 3.1) gives that the constraint is inactive in $(T_{0},T)$. This completes the proof. ∎
Let us also recall a result on vector-valued measures (cf., e.g., page 380 of [13]).
Theorem 4.3 (Dinculeanu-Zinger).
Let $\omega$ be a compact Hausdorff space and let $X$ be a Banach space satisfying the Radon-Nikodym property. Let $\mathcal{F}$ be the collection of Borel sets of $\omega$.
There exists an isomorphism between $(\mathcal{C}^{0}(\omega;X))^{\ast}$ and the space of the regular Borel measures with finite variation taking values in $X^{\ast}$. In particular, for each $F\in(\mathcal{C}^{0}(\omega;X))^{\ast}$, there exists a unique regular Borel measure $\mu:\mathcal{F}\to X^{\ast}$ in $\mathcal{M}(\omega;X^{\ast})$ with finite variation such that
$\langle\langle\alpha,F\rangle\rangle_{X^{\ast},X}=\int_{\omega}\langle\,\mathrm{d}\mu(t),\alpha(t)\rangle_{X^{\ast},X}\,\mathrm{d}t,\quad\textup{ for all }\alpha\in\mathcal{C}^{0}(\omega;X).$
The next lemma aims to verify that the limit $\bm{u}$ introduced in Lemma 4.4 satisfies the initial condition for the velocity. This property hinges on the fact that the confinement condition we are here employing is of Signorini type.
Lemma 4.5.
The limit $\bm{u}$ recovered in (17) satisfies the following properties:
  • $(i)$
    $\dot{\bm{u}}(0)=\bm{u}_{1}$;
  • $(ii)$
    $\ddot{\bm{u}}\in\mathcal{M}([0,T];\bm{V}^{\ast}(\Omega))$.
Proof.
For a.a. $t\in(0,T)$, and for all $\kappa>0$, any solution $\bm{u}_{\kappa}$ for Problem $\mathcal{P}_{\kappa}(\Omega)$ satisfies the following equation:
(18) $\ddot{\bm{u}}_{\kappa}(t)+(\tilde{\mathcal{A}}\bm{u}_{\kappa})(t)+(\tilde{\mathcal{B}}\dot{\bm{u}}_{\kappa})(t)+\dfrac{\chi_{(0,T_{0})}(t)}{\kappa}(\tilde{\mathcal{N}}\bm{u}_{\kappa})(t)=\bm{f}(t),\quad\textup{ in }\bm{V}^{\ast}(\Omega).$
Since $\bm{V}^{\ast}(\Omega)\hookrightarrow\bm{H}^{-1}(\Omega)$, we can test (18) along any element $\bm{v}\in\bm{H}^{1}_{0}(\Omega)$. Thanks to (1), we obtain that the penalty term vanishes if, for a.a. $t\in(0,T)$, the linear operator $\mathcal{N}\bm{u}_{\kappa}(t)$ is tested along an element $\bm{v}\in\bm{H}^{1}_{0}(\Omega)$. Therefore, the function $\bm{u}_{\kappa}$ solves the following equation
(19) $\ddot{\bm{u}}_{\kappa}(t)+(\tilde{\mathcal{A}}\bm{u}_{\kappa})(t)+(\tilde{\mathcal{B}}\dot{\bm{u}}_{\kappa})(t)=\bm{f}(t),\quad\textup{ in }\bm{H}^{-1}(\Omega),$
in the sense of distributions in $(0,T)$. Note that (19) is a linear equation; this implies that:
$\{\ddot{\bm{u}}_{\kappa}\}_{\kappa>0}\textup{ is bounded in }L^{2}(0,T;\bm{H}^{-1}(\Omega))\textup{ independently of }\kappa.$
An application of Theorem 10.1.20 of [34] gives that $\dot{\bm{u}}_{\kappa}\rightharpoonup\dot{\bm{u}}$ in $\mathcal{C}^{0}([0,T];\bm{H}^{-1}(\Omega))$ as $\kappa\to 0$ up to passing to a suitable subsequence, thus verifying the initial condition for the velocity and establishing $(i)$.
To establish that the acceleration $\ddot{\bm{u}}$ enjoys the regularity of a vector-valued measure in the sense of Dinculeanu (cf., e.g., [3, 13]), we first observe that if we test the equation (18) at an arbitrary $\bm{v}\in\bm{V}(\Omega)$ such that $\|\bm{v}\|_{\bm{V}(\Omega)}=1$ and we integrate between 00 and $T$, we obtain that:
(20) $\dfrac{1}{\kappa}\left|\left\langle\int_{0}^{T}\chi_{(0,T_{0})}(t)\mathcal{N}\bm{u}_{\kappa}(t)\,\mathrm{d}t,\bm{v}\right\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\right|$
$\dfrac{1}{\kappa}\left|\int_{0}^{T}\chi_{(0,T_{0})}(t)\langle\mathcal{N}\bm{u}_{\kappa}(t),\bm{v}\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\,\mathrm{d}t\right|$
$\leq\int_{0}^{T}\|\mathcal{A}\bm{u}_{\kappa}(t)\|_{\bm{V}^{\ast}(\Omega)}+\|\mathcal{B}\dot{\bm{u}}_{\kappa}(t)\|_{\bm{V}^{\ast}(\Omega)}\,\mathrm{d}t+\|\dot{\bm{u}}_{\kappa}(T)-\bm{u}_{1}\|_{\bm{V}^{\ast}(\Omega)}.$
We observe that the right-hand side of (20) is bounded independently of $\kappa$ in light of Lemmas 4.24.4, thus showing that:
$\left\{\left\|\dfrac{1}{\kappa}\int_{0}^{T}\chi_{(0,T_{0})}(t)\mathcal{N}\bm{u}_{\kappa}(t)\,\mathrm{d}t\right\|_{\bm{V}^{\ast}(\Omega)}\right\}_{\kappa>0}\textup{ is bounded independently of }\kappa.$
By definition of supremum, for a.a. $t\in(0,T)$, we can find a sequence $\{\bm{v}_{t}^{(k)}\}_{k=1}^{\infty}\subset\bm{V}(\Omega)$ such that $\|\bm{v}_{t}^{(k)}\|_{\bm{V}(\Omega)}=1$, for which,
$\lim_{k\to\infty}\left|\langle\chi_{(0,T_{0})}(t)\mathcal{N}\bm{u}_{\kappa}(t),\bm{v}_{t}^{(k)}\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\right|=\sup_{\begin{subarray}{c}\bm{v}\in\bm{V}(\Omega)\\ \|\bm{v}\|_{\bm{V}(\Omega)=1}\end{subarray}}\left|\langle\chi_{(0,T_{0})}(t)\mathcal{N}\bm{u}_{\kappa}(t),\bm{v}\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\right|,$
and the supremum on the right-hand side of the latter is realised by the vector fields $\bm{v}\in\bm{V}(\Omega)$ such that $\|\bm{v}\|_{\bm{V}(\Omega)}=1$ such that $(v_{i}\bm{e}^{i}\cdot\bm{q})$ has negative sign a.e. in $\Omega$. In correspondence of such elements, we can drop the absolute value. Keeping the latter statement in mind, and applying the Fatou lemma, we obtain that:
$\int_{0}^{T}\left\|\chi_{(0,T_{0})}(t)\mathcal{N}\bm{u}_{\kappa}(t)\right\|_{\bm{V}^{\ast}(\Omega)}\,\mathrm{d}t\leq\liminf_{k\to\infty}\int_{0}^{T}\chi_{(0,T_{0})}(t)\left|\langle\mathcal{N}\bm{u}_{\kappa}(t),\bm{v}_{t}^{(k)}\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\right|\,\mathrm{d}t$
$\leq\sup_{\begin{subarray}{c}\bm{v}\in\bm{V}(\Omega)\\ \|\bm{v}\|_{\bm{V}(\Omega)=1}\end{subarray}}\int_{0}^{T}\chi_{(0,T_{0})}(t)\langle\mathcal{N}\bm{u}_{\kappa}(t),\bm{v}\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\,\mathrm{d}t$
$\leq\sup_{\begin{subarray}{c}\bm{v}\in\bm{V}(\Omega)\\ \|\bm{v}\|_{\bm{V}(\Omega)=1}\end{subarray}}\left|\left\langle\int_{0}^{T}\chi_{(0,T_{0})}(t)\mathcal{N}\bm{u}_{\kappa}(t)\,\mathrm{d}t,\bm{v}\right\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\right|=\left\|\int_{0}^{T}\chi_{(0,T_{0})}(t)\mathcal{N}\bm{u}_{\kappa}(t)\,\mathrm{d}t\right\|_{\bm{V}^{\ast}(\Omega)}.$
In conclusion, we have obtained that
$\int_{0}^{T}\left\|\chi_{(0,T_{0})}(t)\mathcal{N}\bm{u}_{\kappa}(t)\right\|_{\bm{V}^{\ast}(\Omega)}\,\mathrm{d}t=\left\|\int_{0}^{T}\chi_{(0,T_{0})}(t)\mathcal{N}\bm{u}_{\kappa}(t)\,\mathrm{d}t\right\|_{\bm{V}^{\ast}(\Omega)},$
and that:
(21) $\left\{\dfrac{1}{\kappa}\chi_{(0,T_{0})}\tilde{\mathcal{N}}\bm{u}_{\kappa}\right\}_{\kappa>0}\textup{ is bounded in }L^{1}(0,T;\bm{V}^{\ast}(\Omega))\textup{ independently of }\kappa.$
In light of Lemmas 4.24.4 (in particular, here we need the boundedness of the sequence $\{\dot{\bm{u}}_{\kappa}\}_{\kappa>0}$ in $L^{2}(0,T;\bm{V}(\Omega))$ independent of $\kappa$) and (21), we obtain that:
$\{\ddot{\bm{u}}_{\kappa}\}_{\kappa>0}\textup{ is bounded in }L^{1}(0,T;\bm{V}^{\ast}(\Omega))\textup{ independently of }\kappa.$
The latter statement puts us in position to apply Theorem 4.3 so that, up to passing to subsequences, we obtain:
$\ddot{\bm{u}}_{\kappa}\overset{\ast}{\rightharpoonup}\bm{\mu},\quad\textup{ in}\mathcal{M}([0,T];\bm{V}^{\ast}(\Omega))\textup{ as }\kappa\to 0.$
Observe that the vector-valued measure $\bm{\mu}\in\mathcal{M}([0,T];\bm{V}^{\ast}(\Omega))$ can be understood as the acceleration of $\bm{u}$. This interpretation arises from the classical definition of the weak derivative, which states that:
$\int_{0}^{T}\dot{\bm{u}}_{\kappa}(t)\varphi^{\prime}(t)\,\mathrm{d}t=-\int_{0}^{T}\ddot{\bm{u}}_{\kappa}(t)\varphi(t)\,\mathrm{d}t,\quad\textup{ for all }\varphi\in\mathcal{D}(0,T).$
For all $\bm{v}\in\bm{V}(\Omega)$ and all $\varphi\in\mathcal{D}(0,T)$, it results
$\int_{\Omega}\int_{0}^{T}\dot{u}_{i,\kappa}(t)(\varphi^{\prime}(t)v_{i})\,\mathrm{d}t\,\mathrm{d}x=-\int_{\Omega}\int_{0}^{T}\ddot{u}_{i,\kappa}(t)(\varphi(t)v_{i})\,\mathrm{d}t\,\mathrm{d}x,$
so that, letting $\kappa\to 0$ (see Comment 3 of Chapter 4 of [5]) gives
$\int_{\Omega}\left(\int_{0}^{T}\dot{u}_{i}(t)\varphi^{\prime}(t)\,\mathrm{d}t\right)v_{i}\,\mathrm{d}x=\int_{0}^{T}\left(\int_{\Omega}\dot{u}_{i}(t)v_{i}\,\mathrm{d}x\right)\varphi^{\prime}(t)\,\mathrm{d}t=-\int_{0}^{T}\int_{\Omega}\,\mathrm{d}\mu_{i}(t)(v_{i}\varphi(t))\,\mathrm{d}t,$
where the first equality holds by Fubini’s theorem (cf., e.g., [29]). To sum up, we have obtained that
$\int_{0}^{T}\int_{\Omega}\dot{u}_{i}(t)v_{i}\,\mathrm{d}x\varphi^{\prime}(t)\,\mathrm{d}t=-\int_{0}^{T}\int_{\Omega}\,\mathrm{d}\mu_{i}(t)v_{i}\varphi(t)\,\mathrm{d}t,$
for all $\varphi\in\mathcal{D}(0,T)$ and all $\bm{v}\in\bm{V}(\Omega)$. We can therefore interpret the vector-valued measure $\bm{\mu}$ as the acceleration $\bm{u}$. This allows us to adopt the notation $\bm{\mu}=\ddot{\bm{u}}$, and the proof of $(ii)$ is thus complete. ∎
We now observe that further compactness can be extracted in light of Lemma 4.4 and Lemma 4.5.
Lemma 4.6.
Consider the sequence $\{\bm{u}_{\kappa}\}_{\kappa>0}$, where each $\bm{u}_{\kappa}$ is a solution of Problem $\mathcal{P}_{\kappa}(\Omega)$. Up to passing to a subsequence, the following convergences hold:
$\dot{\bm{u}}_{\kappa}\to\dot{\bm{u}},\quad\textup{ in }L^{2}(0,T;\bm{L}^{2}(\Omega))\textup{ as }\kappa\to 0.$
Proof.
Recall that in Lemma 4.5, we have shown that $\{\ddot{\bm{u}}_{\kappa}\}_{\kappa>0}$ is bounded in $L^{1}(0,T;\bm{V}^{\ast}(\Omega))$ independently of $\kappa$. Thanks to (8), we are allowed to apply the Aubin-Lions-Simon theorem (cf., e.g., Theorem 8.62 in [29]) gives that there exists a subsequence, still denoted $\{\dot{\bm{u}}_{\kappa}\}_{\kappa>0}$ such that
$\dot{\bm{u}}_{\kappa}\to\dot{\bm{u}},\quad\textup{ in }L^{2}(0,T;\bm{L}^{2}(\Omega))\textup{ as }\kappa\to 0,$
as it had to be proved. ∎
We are now ready to establish the existence of solutions for Problem $\mathcal{P}(\Omega)$.
Theorem 4.4.
Assume that the applied body force $\bm{f}=(f_{i})\in L^{2}(0,T;\bm{L}^{2}(\Omega))$ is admissible in the sense of Definition 3.1. Then, Problem $\mathcal{P}(\Omega)$ admits at least one solution.
Proof.
Let $\bm{v}\in\mathcal{C}^{0}([0,T];\bm{U}(\Omega))$. Let us test equation (18) at $(\bm{v}(t)-\bm{u}_{\kappa}(t))\in\bm{V}(\Omega)$ for all $t\in[0,T]$. In light of Lemma 4.5, it is straightforward to see that:
(22) $\int_{0}^{T}\langle\ddot{\bm{u}}_{\kappa}(t),\bm{v}(t)\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\,\mathrm{d}t\to\langle\langle\ddot{\bm{u}},\bm{v}\rangle\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)},\quad\textup{ as }\kappa\to 0.$
An application of Lemma 4.6 gives:
(23) $\limsup_{\kappa\to 0}\left(-\int_{0}^{T}\langle\ddot{\bm{u}}_{\kappa}(t),\bm{u}_{\kappa}(t)\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\,\mathrm{d}t\right)$
$=-\lim_{\kappa\to 0}\langle\dot{\bm{u}}_{\kappa}(T),\bm{u}_{\kappa}(T)\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}+\int_{\Omega}\bm{u}_{1}\cdot\bm{u}_{0}\,\mathrm{d}x+\limsup_{\kappa\to 0}\|\dot{\bm{u}}_{\kappa}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{2}$
$=-\langle\dot{\bm{u}}(T),\bm{u}(T)\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}+\int_{\Omega}\bm{u}_{1}\cdot\bm{u}_{0}\,\mathrm{d}x+\|\dot{\bm{u}}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{2}.$
Thanks to Lemma 4.4, we infer that the convergence $\bm{u}_{\kappa}(T)\rightharpoonup\bm{u}(T)$ in $\bm{V}^{\ast}(\Omega)$ as $\kappa\to 0$ holds. This implies that:
(24) $\limsup_{\kappa\to 0}\left(-\int_{0}^{T}\langle\mathcal{B}\dot{\bm{u}}_{\kappa}(t),\bm{u}_{\kappa}(t)\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\,\mathrm{d}t\right)\leq\dfrac{1}{2}\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\bm{u}_{0})e_{i\|j}(\bm{u}_{0})\,\mathrm{d}x$
$\quad-\dfrac{1}{2}\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\bm{u}_{\kappa}(T))e_{i\|j}(\bm{u}_{\kappa}(T))\,\mathrm{d}x.$
In a similar way, we obtain that the convergence $\bm{u}_{\kappa}\rightharpoonup\bm{u}$ in $L^{2}(0,T;\bm{V}(\Omega))$ as $\kappa\to 0$ implies that:
(25) $\limsup_{\kappa\to 0}\left(-\int_{0}^{T}\langle\mathcal{A}\bm{u}_{\kappa}(t),\bm{u}_{\kappa}(t)\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}\,\mathrm{d}t\right)\leq-\int_{0}^{T}\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}(t))e_{i\|j}(\bm{u}(t))\,\mathrm{d}x\,\mathrm{d}t.$
For what concerns the penalty term, we observe that
(26) $-\dfrac{1}{\kappa}\int_{0}^{T}\chi_{(0,T_{0})}(t)\int_{\Gamma}\{[\bm{I}+u_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\}^{-}\left([\bm{I}+v_{i}(t)\bm{e}^{i}]\cdot\bm{q}-[\bm{I}+u_{i,\kappa}(t)\bm{e}^{i}]\cdot\bm{q}\right)\,\mathrm{d}\Gamma\,\mathrm{d}t\leq 0,$
for all $\kappa>0$.
Combining (22)–(26), we obtain that the limit $\bm{u}$ satisfies the following hyperbolic variational inequalities:
$2\rho\langle\langle\ddot{\bm{u}},\bm{v}\rangle\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}-2\rho\langle\dot{\bm{u}}(T),\bm{u}(T)\rangle_{\bm{V}^{\ast}(\Omega),\bm{V}(\Omega)}+2\rho\int_{\Omega}\bm{u}_{1}\cdot\bm{u}_{0}\,\mathrm{d}x+2\rho\|\dot{\bm{u}}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{2}$
$\quad+\int_{0}^{T}\int_{\Omega}A^{ijk\ell}e_{k\|\ell}(\bm{u}(t))e_{i\|j}(\bm{v}(t)-\bm{u}(t))\,\mathrm{d}x\,\mathrm{d}t+\int_{0}^{T}\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\dot{\bm{u}}(t))e_{i\|j}(\bm{v}(t))\,\mathrm{d}x\,\mathrm{d}t$
$\quad-\dfrac{1}{2}\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\bm{u}(T))e_{i\|j}(\bm{u}(T))\,\mathrm{d}x\,\mathrm{d}t+\dfrac{1}{2}\int_{\Omega}B^{ijk\ell}e_{k\|\ell}(\bm{u}_{0})e_{i\|j}(\bm{u}_{0})\,\mathrm{d}x\geq\int_{0}^{T}\int_{\Omega}f^{i}(t)(v_{i}(t)-u_{i}(t))\,\mathrm{d}x\,\mathrm{d}t,$
for all $\bm{v}=(v_{i})\in\mathcal{C}^{0}([0,T];\bm{U}(\Omega))$ or, equivalently, the limit $\bm{u}$ is a solution of Problem $\mathcal{P}(\Omega)$. This completes the proof. ∎
Conclusion and final remarks
In this paper, we demonstrated the existence of solutions for a model that describes the motion of a three-dimensional viscoelastic body constrained within a specific half-space. After defining the concept of “admissible” applied body force, we proposed and justified a concept of solution for this model by studying the asymptotic behaviour of the minimisers of the classical energy of three-dimensional linearly viscoelastic bodies with a relaxation term reflecting the degree to which a particular deformation violates the given constraint. We showed that the solutions for the penalized model converge to the solutions of a system of hyperbolic variational inequalities, where the acceleration term exhibits the regularity of a vector-valued measure as defined by Dinculeanu.
Declarations
Authors’ Contribution
Not applicable.
Acknowledgements
Not applicable.
Ethical Approval
Not applicable.
Availability of Supporting Data
Not applicable.
Competing Interests
All authors certify that they have no affiliations with or involvement in any organisation or entity with any competing interests in the subject matter or materials discussed in this manuscript.
Funding
This research was partly supported by the University Development Fund of The Chinese University of Hong Kong, Shenzhen.